use Cramer's rule to find the solution of the systems of linear equation in terms of the parameter K.


5x - Ky = 6

-2x + 2Ky = -3


1
Expert's answer
2021-04-14T14:02:52-0400

Given,

5x - Ky = 6

-2x + 2Ky = -3

Now comparing the above equation with the general equation,

"ax+by=c"

"a_1=5, b_1=-k, c_1=6"

"a_2=-2, b_2=2k, c=-3"

Now, substituting the values,

"x=\frac{b_2c_1-b_1c_2}{b_2a_1-b_1a_2}"


"=\frac{12k-3k}{10k-2k}"


"=\frac{9k}{8k}"


"=\frac{9}{8}"

"y=\frac{a_2c_1-a_1c_2}{a_2b_1-a_1b_2}"

Now, again substituting the values,

"y=\frac{-12+15}{2k-10k}"

"=\frac{3}{-8k}"